# Assessing the Political Fairness of Multilingual LLMs: A Case Study based on a 21-way Multiparallel EuroParl Dataset

### v1

@@ 6.  Discussion
Our first study on 21-EuroParl reveals significant differences in translation quality of political speeches across different EU parties. This result provides a different perspective on the political biases of multilingual LLMs, favoring the majority parties from left, center, and right over outsiders such as European United Left/Nordic Green Left and Non-attached Members. Our results are consistent across model sizes and families. Further analyses are needed to better understand the origin of these differences, possibly originating either from the pretraining or posttraining data (Gehman et al., 2020; Matic et al., 2020; Hovy and Prabhumoye, 2021; Birhane et al., 2021; Ceron et al., 2025). An other, complementary explanation, would relate these differences to differences in topic or style across EU parties. In any case, these results call for extreme caution when using automatic translation systems in political processes, such as multilingual deliberations (Cabrera, 2024). [^14] As was done during the “Conference on the Future of Europe”. Available online: https://wayback.archive-it.org/12090/20230418091815/https:/futureu.europa.eu/. Besides political fairness, we also found, as expected, unfair translation quality across source languages, with English and Spanish being the best translated, while Estonian and Lithuanian are the worst translated. These differences may of course be due to differences in training data but also to the morphological complexity of these languages Marco et al. (2022). A limit of our Borda metric is that it does not take into account the magnitude of the differences between the sBLEU or COMET scores. In future work, we could add ties to the Borda ranking, based on the significance of the difference (Deutsch et al., 2021; Freitag et al., 2023).
While we focused on EU parties, future work may leverage 21-EuroParl to study the biases regarding national parties or countries. A related question is the effect of the original language (in which the speech was uttered), including the well-known translationese effect (Kurokawa et al., 2009; Lembersky et al., 2012; Volansky et al., 2013; Läubli et al., 2018; Toral et al., 2018; Zhang and Toral, 2019). This is challenging as parties are not equally represented in all original languages, so we could not apply the same Borda method. We found that excluding the translationese subset (where the target language is the original language) did not affect the results of Section 5.3.
Finally, while we focused on LLMs, 21-EuroParl may prove useful to study other Machine Translation algorithms, e.g. examplar-based, statistical, learning or decoding methods.
@@ 7.  Limitations
Given the scale of our experiments (23,386 examples times 420 language pairs times 3 models), we had to rely on automatic metrics to assess translation quality. However, our results should be confirmed using human judgments at a smaller scale before taking any actions.
Again, because of the scale of experiments, we could not afford to run the largest models, e.g. the 235B Qwen3 or 405B Llama-3.1. However, our experiment with Qwen3 shows that, although the performance overall improves from 4B to 8B, the “political bias” stays consistent.
@@ 9.  Acknowledgements
This research was funded by BPI-France under the project AI For Democracy - Democratic Commons, one of seven winners of BPI-France’s “Digital Commons for Generative AI” call for projects, conducted as part of the France 2030 investment plan. We thank all the members of the Democratic Commons program at Sorbonne Université, Sciences Po, and make.org for the discussion and feedback on earlier versions of this work, in particular Nazanin Shafiabadi, Jean-Philippe Cointet, Salim Hafid and David Mas. We also acknowledge the contribution of Hal Daumé III who visited our group in the Spring of 2025. Finally, we thank Ziqian Peng for assistance with sentence alignment. This work was performed using HPC resources from GENCI–IDRIS (Grant 2024-AD010615901).
The authors are not affiliated with any political party.
@@ 10.  Bibliographical References
  • Alves et al. (2024) Duarte Miguel Alves, José Pombal, Nuno M Guerreiro, Pedro Henrique Martins, João Alves, Amin Farajian, Ben Peters, Ricardo Rei, Patrick Fernandes, Sweta Agrawal, Pierre Colombo, José G. C. de Souza, and Andre Martins. 2024. Tower: An open multilingual large language model for translation-related tasks. In First Conference on Language Modeling.
@@
  • Bassani and Romelli (2022) Elias Bassani and Luca Romelli. 2022. ranx.fuse: A python library for metasearch. In CIKM, pages 4808–4812. ACM.
@@
  • Boelaert et al. (2025) Julien Boelaert, Samuel Coavoux, Étienne Ollion, Ivaylo Petev, and Patrick Präg. 2025. Machine bias. how do generative language models answer opinion polls?1. Sociological Methods & Research, 54(3):1156–1196.
@@
  • Martins et al. (2024) PH Martins, P Fernandes, J Alves, NM Guerreiro, R Rei, DM Alves, J Pombal, A Farajian, M Faysse, M Klimaszewski, et al. 2024. Eurollm: Multilingual language models for europe (arxiv: 2409.16235). arxiv.
@@
  • Milmo et al. (2023) Dan Milmo et al. 2023. Chatgpt reaches 100 million users two months after launch. The Guardian, 3:1017–1054.
@@ 11.  Language Resource References
Appendix A Results per Language Pair
A.1.  sBLEU
@@ Appendix A Results per Language Pair
We report average sBLEU results per language pair (source languages in rows and target languages in columns) for each model in Tables 8, 11, 11 and 11. Because the results of Llama-3.2-1B (Table 11) and Llama-3.2-3B (Table 11) are quite poor, we excluded them from the rest of the analysis.
@@
A.2.  COMET
We report COMET results per language pair (source languages in rows and target languages in columns) for each model in Tables 14, 14, and 14.
@@ A.2.  COMET
Appendix B Political Fairness per Target Language
We report the per-target-language sBLEU scores of every EU party for all models in Tables 16 and 16.

### v2

@@ 6.  Discussion
Our first study on 21-EuroParl reveals significant differences in translation quality of political speeches across different EU parties. This result provides a different perspective on the political biases of multilingual LLMs, favoring the majority parties from left and right over outsiders such as European United Left/Nordic Green Left and Non-attached Members. Our results are consistent across model sizes and families. Further analyses are needed to better understand the origin of these differences, possibly originating either from the pretraining or posttraining data (Gehman et al., 2020; Matic et al., 2020; Hovy and Prabhumoye, 2021; Birhane et al., 2021; Ceron et al., 2025; Resnik, 2025). Another, complementary explanation, would relate these differences to differences in topic across EU parties. Indeed, we found that both the party and topic variables had independent effects (see Appendix A.1). In any case, these results call for extreme caution when using automatic translation systems in political processes, such as multilingual deliberations (Cabrera, 2024). [^15] As was done during the “Conference on the Future of Europe”. Available online: https://wayback.archive-it.org/12090/20230418091815/https:/futureu.europa.eu/. Besides political fairness, we also found, as expected, unfair translation quality across source languages, with English and Spanish being the best translated, while Estonian and Lithuanian are the worst translated. These differences may of course be due to differences in training data but also to the morphological complexity of these languages Marco et al. (2022). A limit of our Borda metric is that it does not take into account the magnitude of the differences between the sBLEU or COMET scores. In future work, we could add ties to the Borda ranking, based on the significance of the difference (Deutsch et al., 2021; Freitag et al., 2023).
While we focused on EU parties, future work may leverage 21-EuroParl to study the biases regarding national parties or countries. A related question is the effect of the original language (in which the speech was uttered), including the well-known translationese effect (Kurokawa et al., 2009; Lembersky et al., 2012; Volansky et al., 2013; Läubli et al., 2018; Toral et al., 2018; Zhang and Toral, 2019). This is challenging as parties are not equally represented in all original languages, so we could not apply the same Borda method. We found that excluding the translationese subset (where the target language is the original language) did not affect the results of Section 5.3. Using a two-way ANOVA, we also found that both the party and original language had independent effects (see Appendix A.2).
Finally, while we focused on LLMs, 21-EuroParl may prove useful to study other Machine Translation algorithms, e.g. exemplar-based, statistical, learning or decoding methods.
@@ 7.  Limitations
Given the scale of our experiments (23,386 examples times 420 language pairs times 5 models), we had to rely on automatic metrics to assess translation quality. However, our results should be confirmed using human judgments at a smaller scale before taking any actions.
Again, because of the scale of experiments, we could not afford to run the largest models, e.g. the 235B Qwen3 or 405B Llama-3.1. However, our experiment with Llama-3.1 shows that, although the performance overall improves from 8B to 70B, the “political bias” stays consistent.
@@ 9.  Acknowledgements
We thank the reviewers for their helpful feedback.
This research was funded by BPI-France under the project AI For Democracy - Democratic Commons, one of seven winners of BPI-France’s “Digital Commons for Generative AI” call for projects, conducted as part of the France 2030 investment plan. We thank all the members of the Democratic Commons program at Sorbonne Université, Sciences Po, and make.org for the discussion and feedback on earlier versions of this work, in particular Nazanin Shafiabadi, Jean-Philippe Cointet, Salim Hafid, Léo Labat and David Mas. We also acknowledge the contribution of Hal Daumé III who visited our group in the spring of 2025. Finally, we thank Ziqian Peng for assistance with sentence alignment. This work was performed using HPC resources from GENCI–IDRIS (Grant 2024-AD010615901).
@@ 10.  Bibliographical References
  • Alves et al. (2024) Duarte Miguel Alves, José Pombal, Nuno M Guerreiro, Pedro Henrique Martins, João Alves, Amin Farajian, Ben Peters, Ricardo Rei, Patrick Fernandes, Sweta Agrawal, Pierre Colombo, José G. C. de Souza, and André F.T. Martins. 2024. Tower: An Open Multilingual Large Language Model for Translation-Related Tasks. In First Conference on Language Modeling.
@@
  • Bassani and Romelli (2022) Elias Bassani and Luca Romelli. 2022. ranx.fuse: A Python Library for Metasearch. In Proc. CIKM, pages 4808–4812. ACM.
@@
  • Boelaert et al. (2025) Julien Boelaert, Samuel Coavoux, Étienne Ollion, Ivaylo Petev, and Patrick Präg. 2025. Machine Bias. How Do Generative Language Models Answer Opinion Polls? Sociological Methods & Research, 54(3):1156–1196.
@@
  • Helwe et al. (2025) Chadi Helwe, Oana Balalau, and Davide Ceolin. 2025. Navigating the Political Compass: Evaluating Multilingual LLMs across Languages and Nationalities. In Findings of the Association for Computational Linguistics: ACL 2025, pages 17179–17204, Vienna, Austria. Association for Computational Linguistics.
@@
  • Labat et al. (2026) Léo Labat, Etienne Ollion, and François Yvon. 2026. Polyglots or multitudes? multilingual llm answers to value-laden multiple-choice questions.
@@
  • Martins et al. (2024) Pedro Henrique Martins, Patrick Fernandes, J Alves, Nuno M Guerreiro, Ricardo Rei, Duarte Miguel Alves, José Pombal, Amin Farajian, Manuel Faysse, Mateusz Klimaszewski, Pierre Colombo, Barry Haddow, José G. C. de Souza, Alexandra Birch, and André F.T Martins. 2024. EuroLLM: Multilingual language models for Europe. arxiv preprint 2409.16235.
@@
  • McLean (2019) I McLean. 2019. Voting, page 121–140. Oxford University Press.
  • Milmo (2023) Dan Milmo. 2023. ChatGPT reaches 100 million users two months after launch. The Guardian, 3:1017–1054.
@@
  • Resnik (2025) Philip Resnik. 2025. Large Language Models Are Biased Because They Are Large Language Models. Computational Linguistics, 51(3):885–906.
@@
  • Shafiabadi and Yvon (2026) Nazanin Shafiabadi and François Yvon. 2026. Biases in translation: Assessing opinion distortion in machine translated texts. In Proceedings of the Fifteenth International Conference on Language Resources and Evaluation (LREC’26).
@@
  • Team (2025) Gemma Team. 2025. Gemma 3 Technical Report.
@@ 11.  Language Resource References
  • Yang et al. (2023) Jinrui Yang, Timothy Baldwin, and Trevor Cohn. 2023. Multi-EuP: The Multilingual European Parliament Dataset for Analysis of Bias in Information Retrieval. In Proceedings of the 3rd Workshop on Multi-lingual Representation Learning (MRL), pages 282–291, Singapore. Association for Computational Linguistics.
Appendix A Confounding Variables
A.1.  Topic
Model Metric Party Topic Party $\mathbf{\times}$ Topic
Gemma-3-4B sBLEU 307 329 340
Qwen3-4B sBLEU 331 265 305
Qwen3-8B sBLEU 318 275 335
Llama-3.1-8B sBLEU 255 292 326
Llama-3.1-70B sBLEU 288 331 375
Gemma-3-4B COMET 405 406 398
Qwen3-4B COMET 417 404 402
Qwen3-8B COMET 419 412 415
Llama-3.1-8B COMET 419 409 403
Llama-3.1-70B COMET 411 417 407
Table 8: Number of language pairs, out of 420, for which a two-way ANOVA for the Party and Topic variables gives a statistically significant result (Fisher’s test, $p<0.01$), for sBLEU and COMET scores (the dependent variables)
To understand the results of Section 5.3, we wish to disentangle here the effect of the topic of the text, e.g., if NGL has worse sBLEU scores than EPP, is it because of a political bias of the LLM, or because NGL discusses different topics than EPP that are more complex or otherwise more difficult to translate? To avoid relying on a topic model, we use the EU Committee metadata, each committee being dedicated to a topic. We then conduct a two-way ANOVA for the Party and Topic variables. The results are reported in Table 8. We find that, while the topic has a significant effect, the party also has an independent effect for most language pairs. The effect is consistent for more language pairs with COMET than sBLEU, for both variables.
A.2.  Original Language
Model Metric Party OL Party $\mathbf{\times}$ OL
Gemma-3-4B sBLEU 254 308 406
Qwen3-4B sBLEU 179 331 408
Qwen3-8B sBLEU 238 318 412
Llama-3.1-8B sBLEU 256 414 207
Llama-3.1-70B sBLEU 287 414 242
Gemma-3-4B COMET 405 409 367
Qwen3-4B COMET 417 418 364
Qwen3-8B COMET 419 419 387
Llama-3.1-8B COMET 419 419 337
Llama-3.1-70B COMET 411 420 365
Table 9: Number of language pairs, out of 420, for which a two-way ANOVA for the Party and Original Language (OL) variables gives a statistically significant result (Fisher’s test, $p<0.01$), for sBLEU and COMET scores (the dependent variables)
We conduct the same analysis to study the effect of the original language (in which the speech was uttered). We assumed that the original language might have an effect because of translationese (Kurokawa et al., 2009; Lembersky et al., 2012; Volansky et al., 2013; Läubli et al., 2018; Toral et al., 2018; Zhang and Toral, 2019). However, we found that excluding the translationese subset (where the target language is the original language) did not affect the results of Section 5.3.
Results of the two-way ANOVA are reported in Table 9. As previously, we find that, while the Original Language has an effect, the party has an independent effect for most language pairs. Again, the effect is consistent across more language pairs with COMET than with sBLEU.
Appendix B Results per Language Pair
B.1.  sBLEU
@@ Appendix B Results per Language Pair
We report average sBLEU results per language pair (source languages in rows and target languages in columns) for each model in Tables 10, 11, 12, 15, 15 and 15. Because the results of Llama-3.2-1B (Table 15) and Llama-3.2-3B (Table 15) are quite poor, we excluded them from the rest of the analysis.
src/tgt bg cs da de el en es et fi fr hu it lt lv nl pl pt ro sk sl sv
bg 22.8 27.8 25.2 10.8 40.7 36.2 11.9 16.1 30.8 14.8 25.9 14.0 15.7 23.6 21.9 29.3 30.9 22.4 20.5 24.8
cs 28.9 25.8 23.9 10.3 35.9 32.8 11.3 15.3 28.3 14.3 24.0 13.1 14.4 22.5 21.2 25.7 27.4 29.6 20.2 22.8
da 29.0 21.2 25.4 12.4 40.0 34.3 11.5 16.1 29.3 14.0 25.0 12.8 14.5 23.4 19.9 26.7 28.1 20.5 18.6 29.0
de 28.7 21.7 27.1 11.7 38.4 34.0 11.4 15.6 28.9 14.2 24.2 12.7 14.0 24.6 19.9 26.7 28.1 21.5 19.1 24.7
el 31.0 21.8 26.9 24.5 41.1 36.6 11.2 15.9 30.4 13.9 25.9 12.9 14.4 23.5 20.6 29.8 31.0 21.2 19.6 24.2
en 38.2 27.2 35.3 31.4 13.7 47.3 13.5 19.2 38.5 16.5 31.9 15.4 17.8 27.7 24.9 37.4 39.8 26.5 23.8 31.8
es 31.2 22.2 27.9 25.5 11.0 43.7 11.7 16.0 33.2 14.3 28.3 13.1 14.9 23.9 21.3 33.7 32.8 21.7 19.6 25.3
et 24.2 17.5 21.4 19.7 9.0 29.3 28.2 15.1 23.9 12.8 20.3 11.8 13.0 19.1 17.3 22.2 22.8 17.0 15.5 19.8
fi 24.1 17.5 22.1 20.0 8.4 30.0 28.6 10.8 24.7 12.8 21.0 11.5 12.7 19.8 16.9 22.0 23.3 16.9 15.4 21.4
fr 29.0 20.6 26.0 23.8 9.8 38.7 37.1 10.7 14.7 13.8 28.2 12.5 14.0 23.4 20.1 30.3 30.2 20.0 18.3 23.4
hu 23.9 17.3 21.4 19.7 9.2 29.7 28.9 9.8 13.4 24.0 20.7 10.9 12.1 19.7 16.9 22.9 22.5 16.6 15.4 19.4
it 27.2 19.6 24.3 22.0 10.8 35.7 34.6 10.4 14.0 30.9 13.1 11.8 13.1 22.3 19.1 28.4 27.8 18.8 16.9 22.2
lt 24.2 17.0 20.7 18.6 8.7 28.3 27.2 10.1 13.6 23.3 12.1 19.5 13.8 18.3 16.8 21.5 21.7 16.5 15.4 18.7
lv 25.7 18.2 22.2 20.0 8.6 31.0 29.2 10.7 14.4 24.8 12.9 20.7 13.0 19.5 17.8 23.1 23.3 17.3 16.1 20.0
nl 25.8 19.0 24.1 23.9 10.4 33.3 31.1 10.2 14.3 27.2 13.3 23.0 11.6 12.9 18.3 24.9 25.1 18.1 17.2 22.4
pl 27.4 20.9 23.8 21.6 10.5 33.3 31.5 10.8 14.1 26.9 13.4 22.8 12.5 13.6 21.2 25.3 26.0 20.4 18.1 21.6
pt 30.5 21.4 26.7 24.4 11.3 42.1 40.2 11.1 15.4 32.9 14.0 27.7 12.6 14.2 23.4 20.8 31.4 20.7 18.6 24.3
ro 32.1 22.4 28.3 26.1 10.3 43.7 39.7 11.5 16.0 33.2 14.5 27.9 13.3 14.8 24.1 21.6 32.5 21.7 19.8 25.4
sk 29.9 29.6 26.3 24.5 10.3 37.0 33.8 11.4 15.7 28.8 14.5 24.6 13.2 14.6 22.7 21.5 26.7 28.1 21.1 23.6
sl 28.7 22.8 25.5 23.6 11.1 35.8 33.0 11.0 15.1 28.0 14.0 23.5 12.9 14.2 22.2 20.2 26.1 27.3 22.8 22.7
sv 27.6 19.9 30.6 24.0 9.2 37.3 32.2 11.2 16.2 28.2 13.3 23.9 12.0 13.9 22.5 19.1 25.2 27.0 19.1 17.8
Table 10: sBLEU scores of Gemma-3-4B for each language pair (source language in rows and target language in columns)
@@
src/tgt bg cs da de el en es et fi fr hu it lt lv nl pl pt ro sk sl sv
Table 12: sBLEU scores of Qwen3-8B for each language pair (source language in rows and target language in columns)
@@
B.2.  COMET
We report COMET results per language pair (source languages in rows and target languages in columns) for each model in Tables 18, 18, 18, 20, and 20.
src/tgt bg cs da de el en es et fi fr hu it lt lv nl pl pt ro sk sl sv
bg 91.9 90.3 87.2 70.9 89.5 88.7 82.0 90.1 88.0 84.7 90.0 84.6 84.0 87.9 91.7 88.5 91.8 90.0 87.7 89.9
cs 91.5 90.1 88.4 70.7 89.0 88.1 81.3 90.0 87.6 84.7 89.7 83.2 82.4 88.6 92.3 87.5 91.1 93.2 87.3 90.1
da 91.6 91.6 87.9 74.7 90.0 88.4 81.7 90.6 87.4 84.3 89.7 83.5 82.8 88.2 91.6 88.1 91.3 89.1 86.2 91.6
de 91.1 92.0 90.2 72.7 89.2 87.5 81.3 90.0 87.2 84.1 89.3 82.7 82.2 88.7 91.7 87.4 90.9 90.0 86.4 90.0
el 91.5 91.4 90.0 86.7 89.9 88.7 81.0 89.9 88.2 83.6 90.0 83.9 82.8 87.4 91.4 88.4 91.5 89.7 86.6 89.8
en 93.0 92.5 91.8 89.2 73.6 90.9 84.8 92.6 90.3 86.4 91.5 84.7 84.6 89.7 92.2 90.3 93.4 90.7 87.9 91.6
es 91.7 91.2 89.9 86.6 70.4 89.9 81.6 90.3 89.0 84.5 90.9 83.0 82.2 87.5 91.7 90.1 91.9 89.4 86.3 89.6
et 90.2 90.3 89.0 86.3 69.6 89.2 87.0 89.4 86.8 82.7 88.7 82.2 81.1 87.2 91.0 87.0 89.8 87.2 84.4 88.9
fi 91.1 91.3 90.2 87.1 70.2 90.3 88.4 80.3 88.1 83.1 89.7 82.9 82.0 88.1 91.5 88.0 90.6 88.5 85.2 90.4
fr 91.4 91.0 89.4 86.7 68.9 89.4 89.4 80.7 90.0 83.9 90.9 82.2 82.1 87.8 91.4 89.0 91.6 89.0 85.8 89.0
hu 90.5 90.4 89.0 86.4 69.3 88.8 87.5 79.3 88.5 86.8 89.0 81.2 80.8 87.1 90.8 87.4 90.0 87.7 84.3 88.6
it 91.3 91.1 89.4 86.4 71.4 88.8 89.1 80.3 89.5 88.7 83.4 82.5 81.8 86.9 91.5 88.9 91.4 88.8 85.8 89.2
lt 90.2 89.3 87.9 85.2 69.4 87.3 86.1 78.2 88.1 84.9 82.0 87.8 82.2 85.9 90.4 85.6 88.2 86.1 83.7 87.7
lv 90.8 89.6 88.6 85.9 67.2 88.2 86.4 79.2 88.5 86.0 82.7 88.2 83.4 86.6 90.4 86.1 89.3 86.6 83.7 88.3
nl 90.7 91.0 89.4 87.7 71.1 88.6 87.2 79.9 89.4 87.0 83.4 88.9 82.1 81.5 91.0 86.9 90.4 88.7 85.3 89.2
pl 91.4 92.1 89.5 87.3 71.7 87.8 88.0 80.9 89.8 87.3 84.0 89.6 83.1 82.2 87.8 87.6 91.0 90.1 87.3 89.4
pt 91.6 91.1 89.8 86.6 71.1 89.6 90.3 80.9 90.0 89.0 84.1 90.8 82.4 81.7 87.4 91.4 91.6 88.9 85.8 89.6
ro 92.0 91.6 90.3 87.5 67.5 90.5 89.6 80.1 89.8 88.9 83.7 90.8 82.7 82.0 88.3 91.8 89.3 89.4 86.4 89.9
sk 91.7 94.5 90.2 88.7 69.4 89.2 88.4 81.0 89.9 87.8 84.9 89.9 82.7 82.0 88.8 92.2 87.8 91.3 87.5 90.1
sl 91.4 91.5 89.8 87.6 71.3 88.9 87.9 79.7 89.2 86.8 83.6 89.3 82.5 81.7 87.7 91.7 87.4 90.8 89.4 89.6
sv 91.5 91.7 91.8 87.9 69.6 89.8 88.3 82.0 90.7 87.2 84.1 89.7 83.3 82.6 88.2 91.7 87.9 91.1 89.0 86.3
Table 16: COMET scores of Gemma-3-4B for each language pair (source language in rows and target language in columns)
@@ B.2.  COMET
src/tgt bg cs da de el en es et fi fr hu it lt lv nl pl pt ro sk sl sv
bg 93.2 90.7 88.0 90.7 89.8 89.1 89.7 90.5 88.5 90.3 90.3 90.0 88.8 88.7 92.6 89.1 92.3 91.2 91.3 90.9
cs 90.9 90.9 89.4 90.3 89.7 89.0 90.6 91.6 88.5 90.8 90.3 89.2 87.4 89.9 93.3 88.7 92.1 94.0 88.1 91.5
da 91.0 93.2 88.6 89.7 90.3 88.9 90.6 92.1 88.2 90.6 90.2 89.5 87.9 89.5 92.1 89.0 92.2 91.0 90.7 92.5
de 85.2 92.8 90.8 87.0 89.8 88.3 89.9 90.9 88.1 89.5 89.9 89.0 86.6 89.8 92.6 88.2 91.7 90.9 90.7 91.2
el 73.1 67.5 79.6 79.9 89.7 86.6 72.1 80.5 86.3 70.5 87.4 73.7 68.6 74.6 75.0 82.2 85.9 71.2 83.6 81.9
en 92.8 94.2 92.4 90.2 92.7 91.1 92.9 93.9 90.7 92.5 91.8 91.3 91.1 90.7 93.3 90.8 94.1 92.7 92.5 92.8
es 87.1 89.9 90.5 74.4 89.3 90.3 90.1 91.4 88.6 87.8 90.6 89.3 86.5 88.0 92.3 90.4 92.4 90.5 90.0 90.8
et 90.2 93.0 90.5 88.2 89.1 90.0 88.4 91.6 88.2 90.2 89.9 89.0 88.1 89.1 92.6 88.4 91.4 90.7 89.7 90.9
fi 90.3 93.3 91.3 88.6 90.8 90.9 89.4 91.0 89.2 90.5 90.4 88.9 86.8 89.6 92.9 89.3 91.8 90.9 90.1 92.0
fr 86.7 90.9 89.9 87.1 89.0 89.8 89.7 84.9 90.2 89.8 91.0 85.9 83.6 88.7 92.1 89.4 92.1 89.4 89.7 90.3
hu 90.4 92.4 90.6 87.9 89.1 89.7 88.8 89.5 90.8 88.3 90.0 88.6 87.6 88.8 91.9 88.8 91.1 90.5 89.6 90.6
it 84.7 89.6 89.7 81.8 87.9 89.2 89.6 89.0 83.4 89.3 88.2 86.1 85.1 88.3 91.7 89.4 91.8 90.2 88.3 90.2
lt 89.0 91.0 89.4 86.5 87.3 88.0 87.3 89.3 88.4 86.2 88.1 88.8 88.4 87.0 91.6 86.6 90.1 89.2 88.9 89.4
lv 88.8 91.9 90.0 87.6 88.7 88.8 87.6 90.5 89.0 87.1 89.7 89.3 90.7 88.3 92.0 87.4 90.8 90.2 89.7 90.4
nl 89.4 92.4 90.2 87.6 88.7 89.1 88.0 89.7 91.2 87.9 89.6 89.4 88.5 86.5 91.7 87.8 91.2 90.4 89.5 90.7
pl 90.4 93.2 90.3 84.3 90.2 88.4 88.7 90.2 91.4 88.0 90.3 90.1 89.9 88.2 88.7 88.5 91.7 91.7 91.2 90.9
pt 91.2 92.8 90.8 84.6 91.0 90.1 90.6 90.2 91.6 89.5 90.5 91.2 89.3 87.3 88.7 92.5 92.4 90.9 90.3 91.1
ro 90.7 92.5 91.0 83.5 91.1 91.0 90.1 89.9 89.6 89.4 90.4 91.1 88.4 88.0 89.1 91.7 89.8 90.7 90.0 91.4
sk 90.9 94.6 90.9 89.7 89.2 89.8 89.1 88.2 89.6 88.6 88.6 90.3 87.8 87.5 89.4 93.4 88.8 91.9 90.2 90.6
sl 90.4 92.2 90.8 88.8 88.3 89.6 88.9 88.3 88.8 88.1 88.9 90.3 89.2 86.6 88.1 93.1 88.5 91.8 90.5 90.9
sv 91.0 93.0 92.3 88.4 90.1 90.3 89.0 90.6 91.3 88.1 90.1 90.3 89.6 88.2 89.5 92.6 89.0 92.0 91.0 90.6
Table 20: COMET scores of Llama-3.1-70B for each language pair (source language in rows and target language in columns)
Appendix C Political Fairness per Target Language
We report the per-target-language sBLEU scores of every EU party for all models in Tables 21, 22, 23, and 24.
tgt ALDE ECR EFA EFD EPP NA NGL S&D
bg 27.67 28.70 27.98 28.61 28.67 26.25 27.10 28.72
cs 21.18 21.23 20.63 21.14 21.11 20.47 19.75 21.30
da 26.50 26.12 25.72 25.20 25.76 24.46 24.64 25.88
de 24.02 24.58 22.57 23.46 23.47 20.16 22.29 23.78
el 10.46 10.02 10.64 10.57 10.28 9.92 10.50 10.52
en 36.32 35.88 34.76 36.39 36.42 35.16 35.16 36.71
es 33.16 32.53 31.93 34.04 34.50 32.76 32.96 33.80
et 11.56 11.67 11.72 10.84 11.04 11.04 10.32 11.12
fi 16.10 15.73 16.17 14.67 15.23 15.30 14.66 15.28
fr 28.20 28.17 27.44 28.89 29.38 26.83 28.12 28.92
hu 14.06 13.40 14.43 13.46 14.01 13.80 12.84 13.77
it 24.42 23.89 23.96 24.34 24.76 22.86 23.96 24.54
lt 13.15 13.00 12.94 12.61 12.57 12.78 11.65 12.84
lv 14.43 13.84 14.75 14.17 14.17 13.65 13.16 14.19
nl 23.42 22.85 22.02 21.92 22.20 21.52 20.86 22.94
pl 20.01 19.27 19.19 19.36 19.95 19.42 18.94 20.09
pt 27.33 26.58 26.55 27.26 27.21 26.80 24.78 27.23
ro 27.81 27.77 27.73 27.51 27.72 27.43 26.93 28.06
sk 20.53 21.11 20.18 20.21 20.59 19.80 19.44 20.59
sl 18.47 18.76 18.20 17.67 18.50 17.57 17.55 18.53
sv 24.17 24.21 22.89 23.32 23.39 22.72 22.09 23.43
Table 21: sBLEU scores of Gemma-3-4B per target language. Each language has its performance variance significantly explained by the party variable according to a Kruskal-Wallis test ($p<0.01$).